FIND THE MISSING LENGTH.
LEG A: X
LEG B: X
HYPOTENUSE: 6 times the square root of 2

Answer by checkley77(12569)

X^2+X^2=6SQRT2
2X^2=6*1.414
2X^2=8.485
X^2=8.485/2
X^2=4.2426
X=SQRT4.2426
X=2.06 ANS. FOR THE LENGTH OF THE 2 SIDES.